# A robust parameterized enhanced shift-splitting preconditioner for three-by-three block saddle point problems

### v2

@@ 3 Convergence analysis of the PESS iterative process
[9] Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices. Then, the saddle point matrix $\mathcal{A}$ is positive stable.
@@
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices. Then $\Sigma^{-1}\mathcal{A}$ ((((or, $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})$ is positive stable.
@@
Since the matrices $\Lambda_{1},\Lambda_{2},$ and $\Lambda_{3}$ are SPD, then $\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}$ is SPD, and further $\Lambda_{2}^{-\frac{1}{2}}B\Lambda_{1}^{-\frac{1}{2}}$ and $\Lambda_{3}^{-\frac{1}{2}}C\Lambda_{2}^{-\frac{1}{2}}$ are of full row rank. Consequently, the block structure of $\mathcal{A}$ and the matrix $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}$ are identical. Then by Lemma 3.1, $\mathcal{A}$ is positive stable implies that, the matrix $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}$ and consequently, $\Sigma^{-1}\mathcal{A}$ is positive stable. $\blacksquare$
By [22], the stationary iterative process (2.2) converges to the exact solution of the three-by-three block SPP (1.1) for any initial guess vector if and only if $|\vartheta(\mathcal{T})|<1,$ where $\mathcal{T}$ is the iteration matrix. Now, we establish an if and only if a condition that precisely determines the convergence of the PESS iterative process.
@@
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices, and $s>0$. Then, the PESS iterative process (2.2) converges to the unique solution of the three-by-three block SPP (1.1) if and only if
@@
Since $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}$ is nonsingular then $\mu\neq 0,$ and we have
@@
Using the condition in Theorem 3.1, next, we present two sufficient conditions that ensure the convergence of the PESS iterative process. The first one is presented as follows:
@@
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices, and if $s\geq\frac{1}{2},$ then $\vartheta(\mathcal{T})<1,$ i.e., the PESS iterative process always converges to the unique solution of the three-by-three block SPP (1.1).
@@
By applying Corollary 3.1, the unconditional convergence of the existing preconditioners, namely ${\mathscr{P}}_{SS}$, ${\mathscr{P}}_{GSS},$ and ${\mathscr{P}}_{EGSS}$ can be obtained from ${\mathscr{P}}_{PESS}$ by sustituting $s=\frac{1}{2}$.
@@
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices, if
$s>\max\left\{\frac{1}{2}\left(1-\frac{\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})}{\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}}\right),0\right\},$
then the PESS iterative process is convergent for any random initial guess.
@@
$\Re(\mu)=\frac{{\bf p}^{*}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}}){\bf p}}{2{\bf p}^{*}{\bf p}}\geq\frac{1}{2}\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}}).$
Again, we have $|\mu|\leq\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}),$ then we get the following
$\frac{1}{2}-\frac{\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})}{2\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}}\geq\frac{1}{2}-\frac{\Re(\mu)}{|\mu|^{2}}.$ (3.6)
@@
The sufficient condition on Corollary 3.2 is difficult to find for large $\mathcal{A}$ due to the involvement of computation of $\mu_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})$ and $\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}.$ Thus, we consider $s\geq\frac{1}{2}$ for practical implementation.
@@ 4 Spectral distribution of the PESS preconditioned matrix
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ are full row rank matrices. Suppose that $\lambda$ is an eigenvalue of the PESS preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$. Then, for $s\geq\frac{1}{2},$ we have
@@
Assume that $\lambda$ and $\theta$ are the eigenvalue of the PESS preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and the iteration matrix $\mathcal{T}$, respectively. On the other hand, from the matrix splitting (2.1), we have
@@
Then, we get $\lambda=1-\theta.$ Now, from Corollary 3.1, for $s\geq\frac{1}{2},$ it holds that $|\theta|<1.$ Hence, it follows that, for $s\geq\frac{1}{2},$ $|\lambda-1|<1,$ which completes the proof. $\blacksquare$
@@
Notice that from (4.2a)-(4.2c), we get $\lambda\neq\frac{1}{s},$ otherwise , ${\bf p}=[u^{T},v^{T},w^{T}]^{T}={0}\in{\mathbb{R}}^{n+m+p},$ which is impossible as ${\bf p}$ is an eigenvector.
@@
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices, and $s>0.$ Let $(\lambda,{\bf p}=[u^{T},v^{T},w^{T}]^{T})$ be an eigenpair of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}.$ Then, the following holds
@@ 3. $(3)$
$(1)$ Let $u=0$. Then by (4.2a), we get $(1-s\lambda)B^{T}v=0,$ which shows that $v=0,$ as $B$ has full row rank and $\lambda\neq\frac{1}{s}.$ Furthermore, when combined with (4.2c) leads to $w=0,$ as $\Lambda_{3}$ is SPD. Combining the above facts, it follows that ${\bf p}=[u^{T},v^{T},w^{T}]^{T}=0,$ which is impossible as ${\bf p}$ is an eigenvector. Hence, $u\neq 0.$
$(2)$ Let ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ be an eigenvector corresponding to the real eigenvalue $\lambda.$ Then, we have ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}{\bf p}=\lambda{\bf p},$ or, $\mathcal{A}{\bf p}=\lambda{\mathscr{P}}_{PESS}{\bf p}.$ Consequently, $\lambda$ can be written as
@@
The last inequality follows from the assumptions that $A,\Lambda_{1},\Lambda_{2}$, and $\Lambda_{3}$ are SPD matrices.
@@
Now, the systems of linear equations in (4.3) can also be written as $\mathcal{A}{\bf p}=\mu\Sigma{\bf p},$ or, $\Sigma^{-1}\mathcal{A}{\bf p}=\mu{\bf p},$ where $\mu=\lambda/(1-s\lambda).$ By Lemma (3.2), we have $\Sigma^{-1}\mathcal{A}$ is positive stable, which implies $\Re(\mu)>0,$ hence the proof is completed. $\blacksquare$
@@
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices, and $s>0.$ Suppose $\lambda$ is a real eigenvalue of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}.$ Then
@@
$Au=\lambda(\Lambda_{1}+sA)u,\quad\lambda\Lambda_{3}w=0,$ (4.7)
respectively. The second equation of (4.7) gives $w=0,$ as $\Lambda_{3}$ is SPD. Multiplying $u^{T}$ to the first equation of (4.7) from the left, we get
@@
where $p={\displaystyle{\frac{u^{T}Au}{u^{T}\Lambda_{1}u}}}>0.$ Now, for any nonzero vector $u\in{\mathbb{R}}^{n},$ we can have
@@
where $q={\displaystyle{\frac{u^{T}B^{T}\Lambda_{2}^{-1}Bu}{u^{T}\Lambda_{1}u}}}.$ By solving (4.13), we obtain the following real solutions for $\lambda$
$\lambda^{\pm}=\frac{2sq+p\pm\sqrt{p^{2}-4q}}{2(s^{2}q+sp+1)}$ (4.14)
@@
respectively, where $p,s>0,$ $q\geq 0$ and $p^{2}>4q$. Then $\Phi_{1}$ and $\Phi_{2}$ is monotonically increasing in the argument $p>0,$ and $q\geq 0,$ and decreasing in the argument $q\geq 0$ and $p>0,$ respectively. Now, using (4.8), (4.14) and monotonicity of the functions $\Phi_{1}$ and $\Phi_{2},$ we get the following bounds
@@
Combining (4.15) and (4.16), we get the following bounds for $\lambda$
@@
Since by Remark 4.2, we get $01-s\lambda>$ 0 and the above assertion yields $\mu>\xi_{\max}.$ Hence, the matrix $\mu\Lambda_{1}-A$ is nonsingular and (4.2a) gives
@@
Note that, $v\neq 0.$ Premultiplying by $v^{T}$ on the both sides of (4.21) leads to the following identity
@@
On the other hand, $\mu\Lambda_{1}-A>\Lambda_{1}(\mu-\ \lambda_{\max}(\Lambda^{-1}A))I>0,$ and hence a SPD matrix. Consequently, the matrix $B\left(\mu\Lambda_{1}-A\right)^{-1}B^{T}$ is also a SPD matrix and this implies $v^{T}B\left(\mu\Lambda_{1}-A\right)^{-1}B^{T}v>0,$ for all $v\not=0.$ This leads to a contradiction to (4.22) as $v^{T}C^{T}\Lambda_{3}^{-1}Cv\geq 0.$ Hence, the assumption (4.19) is not true and we get
@@
Again, from ($(1)$) of Proposition 4.1, we have $\lambda>0.$ Combining the above together with (4.9) and (4.17), we obtained the desired bounds in (4.6) for $\lambda.$ Hence, the proof is concluded. ∎
@@
Since ${\mathscr{P}}_{SS}$ is a special case of the PESS preconditioner ${\mathscr{P}}_{PESS}$ for $s=\frac{1}{2}$ as discussed in Table 1, the bounds in (4.24) are obtained by substituting $\Lambda_{1}=\frac{1}{2}\alpha I$ and $s=\frac{1}{2}$ in Theorem 4.2. $\blacksquare$
@@
Let ${\mathscr{P}}_{EGSS}$ be defined as in (1.11) with $\alpha>0$ and let $\lambda$ be an real eigenvalue of the EGSS preconditioned matrix ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}.$ Then,
@@
Since ${\mathscr{P}}_{EGSS}$ is a special case of the PESS preconditioner ${\mathscr{P}}_{PESS}$ for $s=\frac{1}{2}$ as discussed in Table 1, and the bounds are in (4.25) are obtained by substituting $\Lambda_{1}=\frac{1}{2}\alpha P$ and $s=\frac{1}{2}$ in Theorem 4.2. $\blacksquare$
@@
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix, and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices. Let $s>0$ and $\lambda$ be a non-real eigenvalue of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0.$ Suppose ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the eigenvector corresponding to $\lambda.$ Then the following holds

### v3

@@ 3 Convergence analysis of the PESS iterative process
[36] Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices. Then, the saddle point matrix $\mathcal{A}$ is positive stable.
@@
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices. Then $\Sigma^{-1}\mathcal{A}$ ((((or $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})$ is positive stable.
@@
Given that the matrices $\Lambda_{1},\Lambda_{2}$ and $\Lambda_{3}$ are SPD, then $\Lambda_{1}^{-\frac{1}{2}}A\Lambda_{1}^{-\frac{1}{2}}$ is SPD. Furthermore, the matrices $\Lambda_{2}^{-\frac{1}{2}}B\Lambda_{1}^{-\frac{1}{2}}$ and $\Lambda_{3}^{-\frac{1}{2}}C\Lambda_{2}^{-\frac{1}{2}}$ are of full row rank. Consequently, the block structure of $\mathcal{A}$ and the matrix $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}$ are identical. Then by Lemma 3.1, $\mathcal{A}$ is positive stable implies that the matrix $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}$ and consequently, $\Sigma^{-1}\mathcal{A}$ is positive stable. $\blacksquare$
By [41], the stationary iterative process (2.2) converges to the exact solution of the three-by-three block SPP (1.1) for any initial guess vector if and only if $|\vartheta(\mathcal{T})|<1,$ where $\mathcal{T}$ is the iteration matrix. Now, we establish an if and only if condition that precisely determines the convergence of the PESS iterative process.
@@
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices and $s>0$. Then, the PESS iterative process (2.2) converges to the unique solution of the three-by-three block SPP (1.1) if and only if
@@
Since $\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}$ is nonsingular then $\mu\neq 0$ and we have
@@
Using the condition in Theorem 3.1, next, we present two sufficient conditions that ensure the convergence of the PESS iterative process. The first one is presented as follows.
@@
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices and if $s\geq\frac{1}{2},$ then $\vartheta(\mathcal{T})<1,$ i.e., the PESS iterative process always converges to the unique solution of the three-by-three block SPP (1.1).
@@
By applying Corollary 3.1, the unconditional convergence of the existing preconditioners, namely ${\mathscr{P}}_{SS}$, ${\mathscr{P}}_{GSS}$ and ${\mathscr{P}}_{EGSS}$ can be obtained from ${\mathscr{P}}_{PESS}$ by sustituting $s=1/2$.
@@
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices, if
$s>\max\left\{\frac{1}{2}\left(1-\frac{\lambda_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})}{\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}}\right),0\right\},$
then the PESS iterative process is convergent for any initial guess.
@@
$\Re(\mu)=\frac{{\bf p}^{*}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}}){\bf p}}{2{\bf p}^{*}{\bf p}}\geq\frac{1}{2}\lambda_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}}).$
Since $|\mu|\leq\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}}),$ we obtain the following:
$\frac{1}{2}-\frac{\lambda_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})}{2\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}}\geq\frac{1}{2}-\frac{\Re(\mu)}{|\mu|^{2}}.$ (3.6)
@@
The sufficient condition on Corollary 3.2 is difficult to find for large $\mathcal{A}$ due to the involvement of computation of $\lambda_{\min}(\Sigma^{-\frac{1}{2}}(\mathcal{A}+\mathcal{A}^{T})\Sigma^{-\frac{1}{2}})$ and $\vartheta(\Sigma^{-\frac{1}{2}}\mathcal{A}\Sigma^{-\frac{1}{2}})^{2}.$ Thus, we consider $s\geq\frac{1}{2}$ for practical implementation.
@@ 4 Spectral distribution of the PESS preconditioned matrix
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ are full row rank matrices. Suppose that $\lambda$ is an eigenvalue of the PESS preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$. Then, for $s\geq\frac{1}{2},$ we have
@@
Assume that $\lambda$ and $\theta$ are the eigenvalues of the PESS preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and the iteration matrix $\mathcal{T}$, respectively. On the other hand, from the matrix splitting (2.1), we have
@@
Then, we get $\lambda=1-\theta.$ Now, from Corollary 3.1, it holds that $|\theta|<1$ for $s\geq\frac{1}{2}.$ Hence, it follows that, $|\lambda-1|<1$ for $s\geq\frac{1}{2},$ which completes the proof. $\blacksquare$
@@
Notice that from (4.2a)-(4.2c), we get $\lambda\neq 1/s,$ otherwise , ${\bf p}=[u^{T},v^{T},w^{T}]^{T}={0}\in{\mathbb{R}}^{n+m+p},$ which is impossible as ${\bf p}$ is an eigenvector.
@@
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices and $s>0.$ Let $(\lambda,{\bf p}=[u^{T},v^{T},w^{T}]^{T})$ be an eigenpair of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}.$ Then, the following holds:
@@ 3. $(3)$
$(1)$ Let $u=0$. Then by (4.2a), we get $(1-s\lambda)B^{T}v=0,$ which shows that $v=0,$ as $B$ has full row rank and $\lambda\neq 1/s.$ Furthermore, when combined with (4.2c) leads to $w=0,$ as $\Lambda_{3}$ is SPD. Combining the above facts, it follows that ${\bf p}=[u^{T},v^{T},w^{T}]^{T}=0,$ which is impossible as ${\bf p}$ is an eigenvector. Hence, $u\neq 0.$
$(2)$ Let ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ be an eigenvector corresponding to the real eigenvalue $\lambda.$ Then, we have ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}{\bf p}=\lambda{\bf p},$ or $\mathcal{A}{\bf p}=\lambda{\mathscr{P}}_{PESS}\,{\bf p}.$ Consequently, $\lambda$ can be written as
@@
The last inequality follows from the assumptions that $A,\Lambda_{1},\Lambda_{2}$ and $\Lambda_{3}$ are SPD matrices.
@@
The system of linear equations in (4.3) can also be expressed as $\mathcal{A}{\bf p}=\mu\Sigma{\bf p},$ or $\Sigma^{-1}\mathcal{A}{\bf p}=\mu{\bf p},$ where $\mu=\lambda/(1-s\lambda).$ By Lemma (3.2), we have $\Sigma^{-1}\mathcal{A}$ is positive stable, which implies $\Re(\mu)>0,$ hence the proof is completed. $\blacksquare$
@@
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices and $s>0.$ Suppose $\lambda$ is a real eigenvalue of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}.$ Then
@@
$Au=\lambda(\Lambda_{1}+sA)u\leavevmode\nobreak\ \leavevmode\nobreak\ {\text{and}}\leavevmode\nobreak\ \leavevmode\nobreak\ \lambda\Lambda_{3}w=0,$ (4.7)
respectively. The second equation in (4.7) gives $w=0,$ as $\Lambda_{3}$ is SPD. Premultiplying by $u^{T}$ to the first equation of (4.7) , we get
@@
where $p={\displaystyle{\frac{u^{T}Au}{u^{T}\Lambda_{1}u}}}>0.$ Now, for any nonzero vector $u\in{\mathbb{R}}^{n},$ we have
@@
where $q={\displaystyle{\frac{u^{T}B^{T}\Lambda_{2}^{-1}Bu}{u^{T}\Lambda_{1}u}}}.$ By solving (4.13), we obtain the following real solutions for $\lambda$:
$\lambda^{\pm}=\frac{2sq+p\pm\sqrt{p^{2}-4q}}{2(s^{2}q+sp+1)},$ (4.14)
@@
respectively, where $p,s>0,$ $q\geq 0$ and $p^{2}>4q$. Then $\Phi_{1}$ and $\Phi_{2}$ are strictly monotonically increasing in the argument $p>0$ and $q\geq 0,$ and decreasing in the argument $q\geq 0$ and $p>0,$ respectively. Now, using (4.8), (4.14) and monotonicity of the functions $\Phi_{1}$ and $\Phi_{2},$ we get the following bounds:
@@
Combining (4.15) and (4.16), we get the following bounds for $\lambda$:
@@
By Remark 4.2, we get $01-s\lambda>$ 0 and the above assertion yields $\mu>\xi_{\max}.$ Hence, the matrix $\mu\Lambda_{1}-A$ is nonsingular and (4.2a) gives
@@
Note that, $v\neq 0.$ Thus, premultiplying by $v^{T}$ on the both sides of (4.21) leads to the following identity:
@@
On the other hand, $\mu\Lambda_{1}-A>\Lambda_{1}(\mu-\ \lambda_{\max}(\Lambda^{-1}A))I>0$ and hence is a SPD matrix. Consequently, the matrix $B\left(\mu\Lambda_{1}-A\right)^{-1}B^{T}$ is also a SPD matrix and this implies $v^{T}B\left(\mu\Lambda_{1}-A\right)^{-1}B^{T}v>0$ for all $v\not=0.$ This leads to a contradiction to (4.22) as $v^{T}C^{T}\Lambda_{3}^{-1}Cv\geq 0.$ Hence, the assumption (4.19) is not true and we get
@@
Again, we have $\lambda>0$ from $(1)$ of Proposition 4.1. Combining the above together with (4.9) and (4.17), we obtain the desired bounds in (4.6) for $\lambda.$ Hence, the proof is concluded. ∎
@@
Since ${\mathscr{P}}_{SS}$ is a special case of the PESS preconditioner ${\mathscr{P}}_{PESS}$ for $s=1/2$ as discussed in Table 1, the bounds in (4.24) are obtained by substituting $\Lambda_{1}=\frac{1}{2}\alpha I$ and $s=1/2$ in Theorem 4.2. $\blacksquare$
@@
Let ${\mathscr{P}}_{EGSS}$ be defined as in (1.13) with $\alpha>0$ and let $\lambda$ be an real eigenvalue of the EGSS preconditioned matrix ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}.$ Then
@@
Since ${\mathscr{P}}_{EGSS}$ is a special case of the PESS preconditioner ${\mathscr{P}}_{PESS}$ for $s=1/2$ as discussed in Table 1 and the bounds in (4.25) are obtained by substituting $\Lambda_{1}=\frac{1}{2}\alpha P$ and $s=1/2$ in Theorem 4.2. $\blacksquare$
@@
Let $A\in{\mathbb{R}}^{n\times n}$ be a SPD matrix and let $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ be full row rank matrices. Let $s>0$ and $\lambda$ be a non-real eigenvalue of the preconditioned matrix ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $\Im(\lambda)\neq 0.$ Suppose ${\bf p}=[u^{T},v^{T},w^{T}]^{T}$ is the eigenvector corresponding to $\lambda.$ Then the following holds: